# OmniHands: Towards Robust 4D Hand Mesh Recovery via A Versatile Transformer

### v3

@@ 3.4 Losses
3.4.1 maxMSE Loss
Following [16], we use the maxMSE Loss to adjust the weights of each joint to maximize the MSE loss. Given joints locations $\mathcal{P^{*}},\mathcal{P}$, the maxMSE loss is formulated as
$maxMSE(\mathcal{P},\mathcal{P^{*}})=\frac{\sum_{i=1}^{N}||\mathcal{P}_{i}-\mathcal{P^{*}}_{i}||^{4}}{\sum_{i=1}^{N}||\mathcal{P}_{i}-\mathcal{P^{*}}_{i}||^{2}}.$ (10)
@@
3.4.2 Single-Hand Losses
[…]
$L_{mano}=||\tilde{\theta}-\theta||^{2}+||\tilde{\beta}-\beta||^{2}$ (11)
With the pose and shape parameters, we can obtain 3-D vertices $V$ and 3-D joint keypoints $J$. With accurate 3-D ground truth $\tilde{V}$ and $\tilde{J}$, we can supervise the output 3-D keypoints with 3-D loss using maxMSE
$L_{3D}=maxMSE(\tilde{V},V)+maxMSE(\tilde{J},J)$ (12)
With ground truth 3-D joint keypoints and camera parameters, we can get 2-D keypoints $\tilde{j}=\tilde{\pi}(\tilde{J})$ on the original image. The predicted 2-D keypoints are calculated with predicted camera parameter $\pi$. The 2-D projection loss is formulated
$L_{2D}=maxMSE(\pi(J),\tilde{j})$ (13)
3.4.3 Inter-Hand Losses
We utilize Inter-Hand losses in order to optimize the finger touching details of two hands. To calculate the relative distance between right hand keypoints and left hand keypoints, we relocate all keypoints to set the palm joint of right hand on the original coordinate. We denote $\phi_{j}(i,k)\in\mathbb{R}^{3}$ as the distance vector from the $i-th$ right hand joint to the $k-th$ left hand joint, and $\phi_{v}(i,k)\in\mathbb{R}^{3}$ as the distance vector of vertices, defined in the same way.
$L_{jrel}=\sum_{i=1}^{N_{j}}\sum_{k=1}^{N_{j}}||\tilde{\phi}_{j}(i,k)-\phi_{j}(i,k)||^{2}$ (14)
$L_{close}=\sum_{i=1}^{N_{v}}\sum_{k=1}^{N_{v}}\begin{cases}||\tilde{\phi}_{v}(i,k)-\phi_{v}(i,k)||^{2},&\text{if }||\tilde{\phi}_{v}(i,k)||^{2}\leq\alpha\\ 0,&\text{if }||\tilde{\phi}_{v}(i,k)||^{2}>\alpha\end{cases}$ (15)
3.4.4 Sequential Consistency Losses
$L_{consis}=\sum_{t=1}^{N_{t}-1}||\theta_{t+1}-\theta_{t}||^{2}$ (16)
where $N_{t}$ is the length of video sequence.
$L_{consis}=\sum_{i=1}^{N_{t}}\sum_{k=1}^{N_{t}}||J^{*}_{i}-J^{*}_{k}||^{2}$ (17)
4 Experiments
4.1 Implementation details
We implement our model based on PyTorch [43]. We use a pretrained Vision Transformer following a huge design (ViT-h) with 32 layers as the backbone to encode image features. The resolution of input image is $256\times 192$ and the size of the feature map is $H=16,W=12,C=1280$. We use transformer encoders and decoders in our feature fusion modules. We train our model for 1,000,000 steps using AdamW optimizer with a linear learning rate decay schedule from an initial learning rate $1\times 10^{-5}$, and the decay rate is 0.7 for every 200,000. The training set is mixed of Interhand2.6m [39], Re:Interhand [41] and Single Hand Datasets, we use random sampling strategy with sampling rate 40%, 40%, 20%. When taking temporal sequence as input, the length of an image sequence is set to 9 with a gap of 5 frames to sample images in a consecutive video sequence.
4.2 Datasets
Interhand2.6M [39]is the first publicly available dataset of sequences of interacting hands with accurate annotations, which is widely used for training and testing in two-hand reconstruction methods. InterHand2.6M in 30 fps is for video-related research and the 5 fps version is a subset of the 30 fps version which removes the redundancy for single image research. In our method, we make use of 30 fps and 5 fps interacting two-hand (IH) data with human and machine (H+M) annotations and filter invalid data following [32]. We ultimately utilize 1.6M images for training and 1.2M images for testing.
Re:InterHand [41] has images with realistic and diverse appearances along with accurate GT 3D interacting hands. It has 738K video-based images in 3rd-person viewpoints. We take the whole dataset for training.
DexYCB [4] contains sequences of single hands interacting with objects. It consists of 1000 sequences with 582K images. Although it’s not a two-hand dataset, we use it as part of the training set and testing set to enhance our model’s performance in scenarios with object occlusion and motion blur.
Static Single Hand Datasets. To enhance the performance in wild videos, we use single-hand datasets that provide 3D hand annotations for training. Following the previous work [45], we use the datasets with a combination of FreiHAND [75], HO3D [21], MTC [59], RHD [74], H2O3D [21], DEX YCB [4], COCO WholeBody [23], Halpe [14] and MPII NZSL [50].
In-the-wild Datasets. We use in-the-wild datasets HIC [55], ARCTIC [13] and RenderIH [31] to evaluate the performance on various scenarios. HIC provides diverse hand-hand interacting and object-hand interacting sequences and contains 3D GT meshes of both hands. ARCTIC is a multi-view sequential dataset of two-hand interacting with daily use objects, involving occlusions and detailed interactive actions. RenderIH provides realistically rendered hands with a variety of poses, textures, backgrounds, and illuminations, which can simulate various scenarios in reality. For all in-the-wild datasets, we only use them for testing to evaluate the performance on unseen scenarios.
4.3 Metrics
We adopt Mean Per Joint Position Error (MPJPE) and Mean Per Vertex Position Error (MPVPE) and Area Under Curve (AUC) scores to measure the accuracy of the pose and shape of the estimated meshes, and Mean Relative-Root Position Error (MRRPE) to measure the accuracy of relative positions of two hands. For mean position error and AUC we get results under both Root-relative analysis (Ra) and Procrustes analysis (Pa) .Following the prior method [32], we perform root joint alignment and skeleton scaling when evaluating MPJPE. In addition, we use acceleration error Accel_E to evaluate the temporal consistency, following VIBE[27].
4.4 Qualitative Results
For monocular inputs, our qualitative comparisons are shown in Figure 3. We demonstrate results of in-the-wild datasets ARCTIC [13] and RenderIH [31], which are unseen in the training phrase. We compare with IntagHand[32] and ACR [64] in camera perspective and an arbitrary perspective to show more details. We show more results on realistic complex cross-hand interacting samples in Figure 4. The results demonstrate the stability of OmniHands in real-world scenarios.
Figure 3: Qualitative Comparison. We compare OmniHands with state-of-the-art methods on in-the-wild datasets ARCTIC[13] and RenderIH[31]. ARCTIC provides two-hand object interacting images, and RenderIH provides interactive hands under different lighting conditions. We have demonstrated the results from both the camera perspective and an arbitrary perspective. The results show that our method outperforms the state-of-the-art methods in complex real-world environments.
Figure 4: Qualitative Results of in-the-wild hard cases. We show our model’s results on complex cross-hand interactions in realistic scenarios to demonstrate its performance. For a detailed perception, views of multiple perspectives are demonstrated.
4.5 Comparison with state-of-the-art methods
To demonstrate the versatility and stability of OmniHands across various scenarios, we conducted comparative analyses with state-of-the-art methods in multiple domains, including temporal two-hand evaluation, frame-based two-hand evaluation, temporal single hand evaluation and in-the-wild dataset evaluation.
Comparison with Temporal Two Hand Methods. In Table 1 we compare the proposed OmniHands with previous temporal methods on Interhand2.6m 30fps [39]. To demonstrate the benefits of using large mixed training datasets for the model, as well as the impact of using temporal features on model accuracy, we trained four groups of models based on whether additional training data and temporal features were used. For the training datasets, we use Interhand2.6m [39], Re:Interhand [41], and Single Hand Datasets mixed up as our training set for a better model generalization. To ensure a fair comparison with other two-hand reconstruction methods, which are trained only on Interhand2.6m, we use the same training set to train our model and the results are also shown marked with †. As previously mentioned, OmniHands can take sequential frames or single frame as input, to figure out the impact of using temporal features, we report and compare the results of OmniHands with single frame and sequential frames as input, denoting single-frame version with §. The results show that our method far outperforms the existing two-handed temporal models, with 23.34%(7.37 mm $vs.$ 9.09 mm) improvement in MPJPE and 24.05% (2.81 mm/$s^{2}$ $vs.$ 3.70 mm/$s^{2}$) improvement in Accel_E. Comparing the results of whether using temporal features, it can be observed that using temporal feature results in better stability between frames and improvement of key points accuracy. The Accel_E improves 18.86% (2.81 mm/$s^{2}$ $vs.$ 3.34 mm/$s^{2}$) and 17.55% (3.02 mm/$s^{2}$ $vs.$ 3.55 mm/$s^{2}$) respectively, and there is also a notable improvement on MPJPE and MPVPE. By changing the training set, we can find that using extra training data can obviously improve the accuracy and stability.
Comparison with Frame-based Two Hand Methods. To compare with state-of-the-art frame-based methods and demonstrate the performance in single frame input scenarios, we conduct experiments at Interhand2.6m 5fps [39] by expanding single image to sequences. As can be observed at Table 2, even though we drop the temporal knowledge, our method still keep high performance and outperforms previous methods that were specially designed for single frame input scenarios, achieving 7.49 mm, 7.72 mm and 24.58 mm in MPJPE, MPVPE and MRRPE. We also report the results of multi-view frame inputs, without any camera parameter, our model achieves 5.93 mm, 6.20 mm and 14.08 mm in MPJPE, MPVPE and MRRPE, which significantly surpasses the accuracy of monocular estimation.
@@ 4.5 Comparison with state-of-the-art methods
HPS [3] 8.14 19.40 19.78
ITH-3D[65] 6.01 14.45 14.74
SeqHand[62] 6.69 18.20 19.46
Zhao et. al.[73] 3.70 13.45 13.91
Intaghand[32] 4.40 9.26 9.71
DIR[47] 8.92 9.09 9.43
@@
Ours §$\S$ 3.34 7.80 7.95
@@
Table 2: Performance comparisons on the Interhand2.6m 5fps dataset. We compare with the most recent methods. We also report the results of multi-view inputs, marked with $\diamond$.
@@
InterWild[38] 11.12 13.01 29.29
IntagHand[32] 8.79 9.03 -
MEMAHand[57] 8.65 8.89 -
ACR[64] 8.41 8.53 -
Zuo et.al[76] 8.34 8.51 -
DIR[47] 7.51 7.72 28.98
Ours 7.49 7.72 24.58
@@
Table 3: Multi-view single hand recovery results on Dex-YCB, comparing with other single hand multi-view methods.
@@
MVP[69] 6.23 4.26 0.69 9.77 8.14 0.53
PE-Mesh-TR[63] 7.49 4.76 0.64 7.41 4.70 0.64
FTL-Mesh-TR[63] 8.66 5.51 0.59 8.75 5.75 0.59
POEM[63] 6.06 3.93 0.68 6.13 4.00 0.70
Multi-view Fit.[63] 7.33 5.29 0.65 7.22 5.19 0.65
@@
[…]
@@
Arctic[13] IntagHand[32] 24.88 12.02 0.57 0.76 26.30 12.52 0.55 0.75
InterWild[38] 30.21 14.75 0.50 0.71 31.49 15.14 0.48 0.69
DIR[47] 26.63 11.97 0.55 0.76 28.33 12.39 0.52 0.75
ours † 14.21 9.22 0.69 0.78 15.12 9.76 0.63 0.81
ours 13.57 7.98 0.74 0.84 14.51 8.75 0.72 0.82
HIC[55] IntagHand[32] 12.01 7.46 0.76 0.85 12.77 7.82 0.74 0.84
InterWild[38] 10.35 6.32 0.79 0.87 10.68 6.63 0.78 0.86
DIR [47] 11.90 7.12 0.76 0.85 12.79 7.55 0.75 0.84
ours † 9.47 5.55 0.81 0.88 10.35 6.07 0.79 0.87
ours 8.66 4.81 0.82 0.90 9.31 5.30 0.81 0.89
RenderIH[31] IntagHand[32] 33.84 15.19 0.47 0.70 35.17 15.34 0.45 0.69
InterWild [38] 15.92 10.54 0.69 0.79 16.50 10.59 0.68 0.78
DIR[47] 29.48 12.93 0.53 0.74 30.93 12.89 0.51 0.74
ours † 17.22 10.91 0.66 0.78 18.01 11.11 0.64 0.77
ours 13.71 9.17 0.72 0.81 14.32 9.48 0.71 0.81
@@
Comparison with Temporal Single Hand Methods. To further exploit the versatility of our method, we conduct experiments on temporal single hand dataset Dex-YCB [4]. We manage our input by flipping an input image as the opposite hand, and using far away bounding boxes to indicate the hands are not interacting. We report the results of Root-Aligned and Procrustes-aligned metrics. It can be observed in Table 5 that our method outperforms all existing single hand temporal methods in all metrics. It is noteworthy that our lead in Root-Aligned metrics is greater than those in Procrustes-aligned metrics (+38.63% $vs.$ +9.00% in MPJPE, +12.70% $vs.$ +0.89% in JAUC), indicating that we can predict accurate shapes and rotation angles without the assistance of ground truth alignment.
@@
Table 5: Quantitative comparisons with other single hand temoral reconstruction methods on the Dex-YCB dataset.
@@
VIBE[27] 16.95 67.5 6.43 87.1
TCMR[8] 16.03 70.1 6.28 87.5
Tu et al. [54] 19.67 62.5 7.27 85.5
MeshGraphormer[35] 16.21 69.1 6.41 87.2
Deformer[16] 13.64 74.0 5.22 89.6

### v4

@@ 3.4. Losses
3.4.1. maxMSE Loss
Following (Fu et al., 2023), we use the maxMSE Loss to adjust the weights of each joint to maximize the MSE loss. Given joints locations $\mathcal{P^{*}},\mathcal{P}$, the maxMSE loss is formulated as
[…]
@@
3.4.2. Single-Hand Losses
@@
With the pose and shape parameters, we can obtain 3D vertices $V$ and 3D joint keypoints $J$. With accurate 3D ground truth $\tilde{V}$ and $\tilde{J}$, we can supervise the output 3D keypoints with 3D loss using maxMSE
(12) $L_{3D}=maxMSE(\tilde{V},V)+maxMSE(\tilde{J},J)$
@@
3.4.3. Inter-Hand Losses
We utilize Inter-Hand losses in order to optimize the finger touching details of two hands. To calculate the relative distance between right hand keypoints and left hand keypoints, we relocate all keypoints to set the palm joint of right hand on the original coordinate. We denote $\phi_{j}(i,k)\in\mathbb{R}^{3}$ as the distance vector from the ithi-th right hand joint to the kthk-th left hand joint, and $\phi_{v}(i,k)\in\mathbb{R}^{3}$ as the distance vector of vertices, defined in the same way.
[…]
3.4.4. Sequential Consistency Losses
[…]
where $N_{t}$ is the length of video sequence. We choose to constrain joint velocity rather than acceleration in the smoothness loss because acceleration involves second-order differences, which empirically yield weaker gradients and often lead to less stable or effective optimization during training.
[…]
Figure 3. Qualitative Comparison. We compare OmniHands with state-of-the-art methods on in-the-wild datasets ARCTIC(Fan et al., 2023) and RenderIH(Li et al., 2023). ARCTIC provides two-hand object interacting images, and RenderIH provides interactive hands under different lighting conditions. We have demonstrated the results from both the camera perspective and an arbitrary perspective. The results show that our method outperforms the state-of-the-art methods in complex real-world environments.
A Figure.
Figure 4. Qualitative Results of in-the-wild hard cases. We show our model’s results on complex cross-hand interactions in realistic scenarios to demonstrate its performance. For a detailed perception, views of multiple perspectives are demonstrated.
A Figure.
4. Experiments
4.1. Implementation details
We implement our model based on PyTorch (paszke2019pytorch). We use a pretrained Vision Transformer following a huge design (ViT-h) with 32 layers as the backbone to encode image features. The resolution of input image is $256\times 192$ and the size of the feature map is $H=16,W=12,C=1280$. We use transformer encoders and decoders in our feature fusion modules. We train our model for 1,000,000 steps using AdamW optimizer with a linear learning rate decay schedule from an initial learning rate $1\times 10^{-5}$, and the decay rate is 0.7 for every 200,000. The training set is mixed of Interhand2.6m (Moon_2020_ECCV_InterHand2.6M), Re:Interhand (Moon et al., 2024) and Single Hand Datasets, we use a random sampling strategy with sampling rate 40%, 40%, 20%. When taking temporal sequence as input, the length of an image sequence is set to 9 with a gap of 5 frames to sample images in a consecutive video sequence.
4.2. Datasets
Interhand2.6M (Moon_2020_ECCV_InterHand2.6M)is the first publicly available dataset of sequences of interacting hands with accurate annotations, which is widely used for training and testing in two-hand reconstruction methods. InterHand2.6M in 30 fps is for video-related research and the 5 fps version is a subset of the 30 fps version which removes the redundancy for single image research. In our method, we make use of 30 fps and 5 fps interacting two-hand (IH) data with human and machine (H+M) annotations and filter invalid data following (Li et al., 2022). We ultimately utilize 1.6M images for training and 1.2M images for testing.
Re:InterHand (Moon et al., 2024) has images with realistic and diverse appearances along with accurate GT 3D interacting hands. It has 738K video-based images in 3rd-person viewpoints. We take the whole dataset for training.
DexYCB (Chao et al., 2021) contains sequences of single hands interacting with objects. It consists of 1000 sequences with 582K images. Although it’s not a two-hand dataset, we use it as part of the training set and testing set to enhance our model’s performance in scenarios with object occlusion and motion blur.
Static Single Hand Datasets. To enhance the performance in wild videos, we use single-hand datasets that provide 3D hand annotations for training. Following the previous work (Pavlakos et al., 2023), we use the datasets with a combination of FreiHAND (Zimmermann et al., 2019), HO3D (Hampali et al., 2020), MTC (Xiang et al., 2019), RHD (Zimmermann and Brox, 2017), H2O3D (Hampali et al., 2020), DEX YCB (Chao et al., 2021), COCO WholeBody (Jin et al., 2020), Halpe (Fang et al., 2022) and MPII NZSL (Simon et al., 2017). We simulate a pair of widely separated hands by randomly combining two single-hand images and assigning them two non-overlapping bounding boxes.
In-the-wild Datasets. We use in-the-wild datasets HIC (Tzionas et al., 2016), ARCTIC (Fan et al., 2023) and RenderIH (Li et al., 2023) to evaluate the performance on various scenarios. HIC provides diverse hand-hand interacting and object-hand interacting sequences and contains 3D GT meshes of both hands. ARCTIC is a multi-view sequential dataset of two-hand interacting with daily use objects, involving occlusions and detailed interactive actions. RenderIH provides realistically rendered hands with a variety of poses, textures, backgrounds, and illuminations, which can simulate various scenarios in reality. For all in-the-wild datasets, we only use them for testing to evaluate the performance on unseen scenarios.
4.3. Metrics
We adopt Mean Per Joint Position Error (MPJPE) and Mean Per Vertex Position Error (MPVPE) and Area Under Curve (AUC) scores to measure the accuracy of the pose and shape of the estimated meshes, and Mean Relative-Root Position Error (MRRPE) to measure the accuracy of relative positions of two hands. For mean position error and AUC we get results under both Root-relative analysis (Ra) and Procrustes analysis (Pa). Following the prior method (Li et al., 2022), we perform root joint alignment and skeleton scaling when evaluating MPJPE. In addition, we use acceleration error Accel_E to evaluate the temporal consistency, following VIBE(Kocabas et al., 2020). To better evaluate the accuracy of two-hand interactions, we propose Mean Inter-Joint Error (MIJE) as an evaluation metric.
4.4. Qualitative Results
For monocular inputs, our qualitative comparisons are shown in Figure 3. We demonstrate results of in-the-wild datasets ARCTIC (Fan et al., 2023) and RenderIH (Li et al., 2023), which are unseen in the training phase. We compare with IntagHand(Li et al., 2022), ACR (Yu et al., 2023) and InterWild(Moon, 2023) in camera perspective and an arbitrary perspective to show more details. We show more results on realistic complex cross-hand interacting samples in Figure 4, comparing with InterWild. The results demonstrate the stability of OmniHands in real-world scenarios and in complicated interaction cases.
4.5. Comparison with state-of-the-art methods
To demonstrate the versatility and stability of OmniHands across various scenarios, we conducted comparative analyses with state-of-the-art methods in multiple domains, including temporal two-hand evaluation, frame-based two-hand evaluation, temporal single hand evaluation, frame-based single hand evaluation and in-the-wild dataset evaluation.
[…]
Comparison with Frame-based Two Hand Methods. To compare with state-of-the-art frame-based methods and demonstrate the performance in single frame input scenarios, we conduct experiments at Interhand2.6m 5fps (Moon_2020_ECCV_InterHand2.6M) by setting the sequence length to 1. As can be observed at Table 2, even though we drop the temporal knowledge, our method still keep high performance and outperforms previous methods that were specially designed for single frame input scenarios, achieving 7.49 mm, 7.72 mm and 24.58 mm in MPJPE, MPVPE and MRRPE. The improvement of our model becomes more significant when trained on the full dataset, achieving 6.77 mm, 6.96 mm and 22.01 mm in MPJPE, MPVPE and MRRPE. We also report the results of multi-view frame inputs, without any camera parameter, our model achieves 5.93 mm, 6.20 mm and 14.08 mm in MPJPE, MPVPE and MRRPE, which significantly surpasses the accuracy of monocular estimation.
Table 1. Quantitative comparisons with frame-based and temporal methods on Interhand2.6m 30fps dataset. The sign $\dagger$} stands for using only Interhand2.6m 30fps as training set, and $\S$ stands for using frame-based input.
[…]
HPS (boukhayma20193d) 8.14 19.40 19.78
ITH-3D(Zhang et al., 2021) 6.01 14.45 14.74
SeqHand(Yang et al., 2020) 6.69 18.20 19.46
Zhao et. al.(Zhao et al., 2024) 3.70 13.45 13.91
Intaghand(Li et al., 2022) 4.40 9.26 9.71
DIR(Ren et al., 2023) 8.92 9.09 9.43
Ours $\dagger$} $\S$ 3.55 7.99 8.15
Ours $\dagger$} 3.02 7.59 7.79
Ours $\S$ 3.34 7.80 7.95
@@ 4.5. Comparison with state-of-the-art methods
[…]
InterWild(Moon, 2023) 14.76 13.24 26.25
IntagHand(Li et al., 2022) 8.79 9.02 27.74
MEMAHand(Wang et al., 2023) 8.65 8.89 -
ACR(Yu et al., 2023) 8.41 8.53 -
Zuo et.al(Zuo et al., 2023) 8.34 8.51 -
DIR(Ren et al., 2023) 7.51 7.72 28.98
[…]
Ours 6.77 6.96 22.01
@@
Table 3. Multi-view single hand recovery results on Dex-YCB, comparing with other single hand multi-view methods. The model reported was trained on Dex-YCB.
@@
MVP(MVPzhang2021direct) 6.23 4.26 0.69 9.77 8.14 0.53
PE-Mesh-TR(yang2023poem) 7.49 4.76 0.64 7.41 4.70 0.64
FTL-Mesh-TR(yang2023poem) 8.66 5.51 0.59 8.75 5.75 0.59
POEM(yang2023poem) 6.06 3.93 0.68 6.13 4.00 0.70
Multi-view Fit.(yang2023poem) 7.33 5.29 0.65 7.22 5.19 0.65
@@
Table 4. Comparison of two-hand reconstruction methods on in-the-wild datasets. Arctic, HIC and RenderIH are taken as test set, which are unseen during training. Our method marked with $\dagger$} is trained on only Interhand2.6m, and the unmarked one is trained on mixed training set.
@@
Arctic(Fan et al., 2023) IntagHand(Li et al., 2022) 24.88 12.02 0.57 0.76 203.04 26.30 12.52 0.55 0.75
InterWild(Moon, 2023) 22.74 11.31 0.59 0.72 143.19 23.92 11.74 0.56 0.71
DIR(Ren et al., 2023) 26.63 11.97 0.55 0.76 253.49 28.33 12.39 0.52 0.75
ours † 14.21 9.22 0.69 0.78 112.81 15.12 9.76 0.63 0.81
ours 13.57 7.98 0.74 0.84 105.43 14.51 8.75 0.72 0.82
HIC(Tzionas et al., 2016) IntagHand(Li et al., 2022) 12.01 7.46 0.76 0.85 36.46 12.77 7.82 0.74 0.84
InterWild(Moon, 2023) 11.23 6.62 0.79 0.86 30.35 11.65 6.99 0.76 0.86
DIR (Ren et al., 2023) 11.90 7.12 0.76 0.85 32.83 12.79 7.55 0.75 0.84
ours † 9.47 5.55 0.81 0.88 26.84 10.35 6.07 0.79 0.87
ours 8.66 4.81 0.82 0.90 26.56 9.31 5.30 0.81 0.89
RenderIH(Li et al., 2023) IntagHand(Li et al., 2022) 33.84 15.19 0.47 0.70 80.28 35.17 15.34 0.45 0.69
InterWild (Moon, 2023) 19.12 13.16 0.62 0.72 56.10 19.74 12.77 0.61 0.73
DIR(Ren et al., 2023) 29.48 12.93 0.53 0.74 78.53 30.93 12.89 0.51 0.74
ours † 17.22 10.91 0.66 0.78 45.94 18.01 11.11 0.64 0.77
ours 13.71 9.17 0.72 0.81 42.26 14.32 9.48 0.71 0.81
Figure 5. The qualitative results of OmniHands with multi-view inputs. We present the results on Interhand2.6m (Moon_2020_ECCV_InterHand2.6M) at row 1 to row 5 and results on Arctic (Fan et al., 2023) at row 6 to row 9. In each row, we present pairs of input and output images from 4 different camera views. The images on the left side of each pair serve as the multi-view sequence input.
A figure.
Comparison on Single hand Dataset. Although our framework is primarily designed for interacting hands reconstruction, it also achieves performance comparable to state-of-the-art single-hand methods on the task of single-hand reconstruction. We test OmniHands on 2 widely used benchmarks: Freihand (Zimmermann et al., 2019) and HO3Dv2 (Hampali et al., 2020), , with the results presented in Table 6, Table 7. For a fair comparison, we adopt the same training settings as Hamba (dong2024hamba). Specifically, the model in Table 6 is trained only using the FreiHAND dataset, and the model in Table 7 is trained on the same datasets as HaMeR (Pavlakos et al., 2023).
In the task of interacting-hand reconstruction, we primarily use MRRPE to evaluate the accuracy of the relative root positions, specifically the wrist positions, between the two hands as an indicator of interaction accuracy. However, such a numerical metric does not effectively capture the front-back ordering of the fingers between the hands, which can lead to noticeable differences in visual quality. Therefore, we propose Mean Inter-Joint Error (MIJE) to evaluate the hand interaction performance. The MIJE is computed as follows:
(18) $\mathbf{MIJE}=\frac{1}{|J_{r}|\cdot|J_{l}|}\sum_{j_{r}\in J_{r}}\sum_{j_{l}\in J_{l}}\|j_{r}-j_{l}\|_{2}$
We evaluated this metric on InterHand2.6M and compared it with state-of-the-art methods, the results are reported in Table 5. As shown, our method achieves notably better performance in handling complex two-hand interaction poses. This advantage is also reflected in other qualitative results.
Table 5. The Mean Inter-Joint Error (MIJE) metric, measured in millimeters, of our method compared to the SOTA in two-hand reconstruction on the InterHand2.6M dataset. Lower values indicate better performance.
Ours IntagHand InterWild DIR
26.56 32.36 58.51 73.40
@@
Comparison with Temporal Single Hand Methods. To further exploit the versatility of our method, we conduct experiments on temporal single hand dataset Dex-YCB (Chao et al., 2021). We manage our input by flipping an input image as the opposite hand, and using far away bounding boxes to indicate the hands are not interacting. We report the results of Root-Aligned and Procrustes-aligned metrics. It can be observed in Table 8 that our method outperforms all existing single hand temporal methods in all metrics. It is noteworthy that our lead in Root-Aligned metrics is greater than those in Procrustes-aligned metrics (+38.63% $vs.$ +9.00% in MPJPE, +12.70% $vs.$ +0.89% in JAUC), indicating that we can predict accurate shapes and rotation angles without the assistance of ground truth alignment.
@@
Table 6. Quantitative comparisons with other single hand reconstruction methods on Freihand dataset.
Method PA-MPJPE $\downarrow$ PA-MPVPE $\downarrow$ F@5mm $\uparrow$ F@15mm $\uparrow$
PointHMR(kim2023pointhmr) 6.1 6.6 0.720 0.984
zhou et. al (zhou2024simple) 5.7 6.0 0.772 0.986
HaMeR (Pavlakos et al., 2023) 6.0 5.7 0.785 0.990
Hamba (dong2024hamba) 5.8 5.5 0.798 0.991
Ours 5.2 5.7 0.779 0.993
Table 7. Quantitative comparisons with other single hand reconstruction methods on HO3Dv2 dataset.
Method PA-MPJPE $\downarrow$ PA-MPVPE $\downarrow$ F@5mm $\uparrow$ F@15mm $\uparrow$
H2ONet (xu2023h2onet) 8.5 8.6 0.966 0.570
HandBooster (xu2024handbooster) 8.2 8.4 0.972 0.585
HaMeR (Pavlakos et al., 2023) 7.6 7.9 0.980 0.635
Hamba (dong2024hamba) 7.5 7.7 0.982 0.648
Ours 6.9 7.7 0.985 0.616
Table 8. Quantitative comparisons with other single hand temporal reconstruction methods on the Dex-YCB dataset. The model reported was trained on Dex-YCB
@@
HaMeR (Pavlakos et al., 2023) 19.45 68.4 6.09 87.8
VIBE(Kocabas et al., 2020) 16.95 67.5 6.43 87.1
TCMR(Choi et al., 2021) 16.03 70.1 6.28 87.5
Tu et al. (Tu et al., 2023) 19.67 62.5 7.27 85.5
MeshGraphormer(Lin et al., 2021) 16.21 69.1 6.41 87.2
Deformer(Fu et al., 2023) 13.64 74.0 5.22 89.6
@@
Quantitative Comparison on In-the-wild Datasets. We report the results on three in-the-wild datasets: Arctic (Fan et al., 2023), HIC (Tzionas et al., 2016) and RenderIH (Li et al., 2023), to compare the performance at in-the-wild scenarios. Arctic contains images of two hand interacting with objects, including a large number of cases with hand-object occlusion and hand-hand occlusion. HIC is a small dataset of two hand interaction sequences. And RenderIH contains images of complicated two hand interactions under various illumination conditions. We only use samples in which both hands are visible. For fair comparison, temporal data or multi-view data are not used in these experiments, a single image is taken as input for each case.
In Table 4 we compare the results with IntagHand (Li et al., 2022), InterWild (Moon, 2023) and DIR (Ren et al., 2023). IntagHand (Li et al., 2022) and DIR (Ren et al., 2023) are two open-source inter-hand reconstruction models that have demonstrated excellent performance in training and testing on Interhand2.6m (Moon_2020_ECCV_InterHand2.6M). InterWild (Moon, 2023) is a SOTA method specifically designed for two-hand recovery in the wild. To demonstrate the efficiency of using large training data, we report results of models trained under two training set. One of our reported models is trained solely on InterHand2.6M, using the same training set as other comparison methods to enable a fair quantitative comparison; this model is marked with $\dagger$}. The other unmarked model is trained on a mixed dataset comprising InterHand2.6M, Re:InterHand, and single hand datasets.